Question 5 0.49/1 View Policies Show Attempt History Current Attempt in Progress The stockholders equity section of Metlock1. No Entry 223510 Discount on Bonds Payable 8680 Bonds Payable 217000 Paid-in Capital-Stock Warrants 15190 262200 Common Sto(b) Prepare the stockholders equity section of the balance sheet at the end of the current year. Assume that retained earnin
Answer #1

A)Number of stock warrants issued = 217000 bonds *1 warrant /100 =2170 warrants

Number of warrants exercised in part 4 = 2170 *80%= 1736

Date Account title Debit credit
4 Paid in capital -stock warrant 12152
cash 48608
common stock (1736*10 par value) 17360
paid in capital in excess of par -common stock (residual value) 43400

B)

Metlock Inc

Balance sheet

Stockholders equity
Paid in capital
common stock 3002960
Paid in capital in excess of par -common stock 1019160
Paid in capital -stock warrants 3038
Total paid in capital 4025158
Retained earning 707000
Total stockholders equity 4732158

#Common stock =2810000 beginning+87400 (part 3 ) +17360 (part 4) +88200 (part 6) = 3002960

Paid in capital in excess of par -common stock= 554000 beginning+174800 (part 3 ) +43400 (part 4) +246960 (part 6) = 1019160

Paid in capital-stock warrant =15190 (part 2) -12152 (part 3)expired =3038